OpenStudy (anonymous):

A spacecraft is in orbit around a planet. The radius of the orbit is 2.9 times the radius of the planet (which is R = 71498 km). The gravitational field at the surface of the planet is 24 N/kg. What is the period of the spacecraft's orbit? [Hint: You don't need any more data about the planet to solve the problem.]

OpenStudy (ash2326):

Radius of the orbit= 2.9 \(\times\) 71498= 207344.2 km We have Gravitational Field at the surface as 24 N/KG Let E be gravitational field \[E= G\frac{M}{R^2}\] G=universal gravitational constant M= mass of planet, Let's find the mass \[M= \frac{E\times R^2}{G}\] We get mass of planet \[M=\frac{24N/kg\times (71498*1000 m)^2}{6.673\times 10^{-11} N-m^2/kg^2}\] we get \[\large{M=1.83\times 10^{27} kg}\] From Kepler's third law time period of a satellite, T \[T=\sqrt{\frac{4\times \pi\times r^3}{G\times M}}\] r= radius of orbit \[T=\sqrt{\frac{4\times\pi\times(207344.2 *1000m)^3}{6.673\times10^{-11} Nm^2/kg^2\times 1.83\times10^{27} kg }}\] Now you can find the period, can't you??

OpenStudy (ash2326):

a slight correctiom \[ T=\sqrt{\frac{4\times\pi^2\times r^3}{G\times M}}\] now substitute the values and find the orbital period

OpenStudy (anonymous):

is that given in Seconds? Because It has to be in hours..and I need to know if I need to convert anything beforehand

OpenStudy (ash2326):

It's given in seconds, divide it by 3600 to get in hours